# MVSplat: Efficient 3D Gaussian Splatting from Sparse Multi-View Images

### v1

@@ Abstract
We propose MVSplat, an efficient feed-forward 3D Gaussian Splatting model learned from sparse multi-view images. To accurately localize the Gaussian centers, we propose to build a cost volume representation via plane sweeping in the 3D space, where the cross-view feature similarities stored in the cost volume can provide valuable geometry cues to the estimation of depth. We learn the Gaussian primitives’ opacities, covariances, and spherical harmonics coefficients jointly with the Gaussian centers while only relying on photometric supervision. We demonstrate the importance of the cost volume representation in learning feed-forward Gaussian Splatting models via extensive experimental evaluations. On the large-scale RealEstate10K and ACID benchmarks, our model achieves state-of-the-art performance with the fastest feed-forward inference speed (22 fps). Compared to the latest state-of-the-art method pixelSplat, our model uses 10×10$\times$10 × fewer parameters and infers more than 2×2$\times$2 × faster while providing higher appearance and geometry quality as well as better cross-dataset generalization.
@@ 1 Introduction
3D scene reconstruction and novel view synthesis from very sparse (i.e., as few as two) images pose a fundamental challenge in computer vision. While remarkable progress has been made using neural scene representations, e.g., Scene Representation Networks (SRN) [25], Neural Radiance Fields (NeRF) [16] and Light Filed Networks (LFN) [24], these methods are still not satisfactory for practical applications due to expensive per-scene optimization [18, 29, 32], high memory cost [2, 11, 33] and slow rendering speed [30, 38].
@@
Subsequently, several feed-forward Gaussian Splatting methods have been proposed to explore 3D reconstruction from sparse view images, notably Splatter Image [27] and pixelSplat [1]. Splatter Image regresses pixel-aligned Gaussian parameters from a single view using a U-Net architecture, which achieves promising results for single object 3D reconstruction. However, 3D reconstruction from a single image is inherently ill-posed and ambiguous, which makes it particularly challenging to apply to more general and larger scene-level reconstruction, the key focus of our paper. For general scene reconstruction, pixelSplat [1] proposes to regress Gaussian parameters from two input views. Specifically, pixelSplat predicts a probabilistic depth distribution from features, and then depths are sampled from the learned distribution. Despite pixelSplat learns cross-view-aware features with an epipolar transformer, it is still challenging to predict a reliable probabilistic depth distribution solely from image features, making pixelSplat’s geometry reconstruction of comparably low quality and exhibit noisy artifacts (see Fig. 1 and Fig. 4). For improved geometry reconstruction results, slow depth finetuning with an additional depth regularization loss is required.
@@
To accurately localize the 3D Gaussian centers, we propose to build a cost volume representation via plane sweeping [37, 35] in 3D space. Specifically, the cost volume stores cross-view feature similarities for all potential depth candidates, where the similarities can provide valuable geometry cues to the localization of 3D surfaces (e.g., high similarity more likely indicates a surface point). With our cost volume representation, the task is formulated as learning to perform feature matching to identify the Gaussian centers, unlike the data-driven 3D regression from image features in previous works [27, 1]. Such a formulation reduces the task’s learning difficulty, enabling our method to achieve state-of-the-art performance with lightweight model size and fast speed.
@@
We obtain 3D Gaussian centers by unprojecting the multi-view-consistent depths estimated by our constructed multi-view cost volumes with a 2D network. Additionally, we also predict other Gaussian properties (covariance, opacity and spherical harmonics coefficients), in parallel with the depths. This enables the rendering of novel view images using the predicted 3D Gaussians with the differentiable splatting operation [12]. Our full model MVSplat is trained end-to-end purely with the photometric loss between rendered and ground truth images.
@@
On the large-scale RealEstate10K [42] and ACID [14] benchmarks, our cost volume-based method MVSplat achieves state-of-the-art performance with the fastest feed-forward inference speed (22 fps). Compared to the state-of-the-art pixelSplat [1] (see Fig. 1), our model uses 10×10$\times$10 × fewer parameters and infers more than 2×2$\times$2 × faster while providing higher appearance and geometry quality as well as better cross-dataset generalization. Extensive ablation studies and analysis underscore the significance of our feature matching-based cost volume design in enabling highly efficient feed-forward 3D Gaussian Splatting models.
@@ 2 Related Work
Sparse View Scene Reconstruction and Synthesis. The original NeRF and 3DGS methods are both designed for very dense views (e.g., 100) as inputs, which can be tedious to capture for real-world applications. Recently, there have been growing interests [18, 29, 2, 3, 33, 1, 32] in scene reconstruction and synthesis from sparse input views (e.g., 2 or 3). Existing sparse view methods can be broadly classified into two categories: per-scene optimization and cross-scene feed-forward inference methods. Per-scene approaches mainly focus on designing effective regularization terms [18, 29, 39, 4, 32] to better constrain the optimization process. However, they are inherently slow at inference time due to the expensive per-scene gradient back-propagation process. In contrast, feed-forward models [38, 2, 3, 33, 1, 27] learn powerful priors from large-scale datasets, so that 3D reconstruction and view synthesis can be achieved via a single feed-forward inference by taking sparse views as inputs, which makes them significantly faster than per-scene optimization methods.
@@
Feed-Forward NeRF. pixelNeRF [38] pioneered the paradigm of predicting pixel-aligned features from images for radiance field reconstruction. The performance of feed-forward NeRF models progressively improved with the use of feature matching information [2, 3], Transformers [21, 6, 17] and 3D volume representations [2, 33]. The state-of-the-art feed-forward NeRF model MuRF [33] is based on a target view frustum volume and a (2+1)D CNN for radiance field reconstruction, where the 3D volume and CNN need to be constructed and inferred for every target view. This makes MuRF expensive to train, with comparably slow rendering. Most importantly, all existing feed-forward NeRF models suffer from an expensive per-pixel volume sampling in the rendering process.
@@
Feed-Forward 3DGS. 3D Gaussian Splatting avoids NeRF’s expensive volume sampling via a rasterization-based splatting approach, where novel views can be rendered very efficiently from a set of 3D Gaussian primitives. Very recently, a few feed-forward 3DGS models have been proposed to solve the sparse view to 3D task. Splatter Image [27] proposes to regress pixel-aligned Gaussian parameters from a single view with a U-Net. However, it mainly focuses on single object reconstruction, while we target a more general setting and larger scenes. pixelSplat [1] proposes to regress Gaussian parameters from two input views, where the epipolar geometry is leveraged to learn cross-view aware features. However, it estimates depth by sampling from a predicted distribution, which is ambiguous and can lead to poor geometry reconstruction. In contrast, we learn to predict depth directly from the feature matching information encoded within a cost volume, which makes it more geometry-aware and leads to a more lightweight model (10×10$\times$10 × fewer parameters and more than 2×2$\times$2 × faster) and significantly better geometries. Another related work GPS-Gaussian [41] proposes a feed-forward Gaussian model for the reconstruction of humans, instead of general scenes. It relies on two rectified stereo images to estimate the disparity, while our method works for general unrectified multi-view posed images. During training, GPS-Gaussian requires ground truth depth for supervision, while our model is trained from RGB images alone.
@@
Figure 2: Overview of MVSplat. Given multiple posed images as input, we first extract multi-view image features with a multi-view Transformer, which contains self- and cross-attention layers to exchange information across views. Next, we construct per-view cost volumes using plane sweeping. The Transformer features and cost volumes are concatenated together as input to a 2D U-Net (with cross-view attention) for cost volume refinement and predicting per-view depth maps. The per-view depth maps are unprojected to 3D and combined using a simple deterministic union operation as the 3D Gaussian centers. The opacity, covariance and color Gaussian parameters are predicted jointly with the depth maps. Finally, novel views are rendered from the predicted 3D Gaussians with the splatting operation.
@@ 3 Method
To enable high-quality rendering and reconstruction, it is crucial to predict the position $\bm{\mu}_{j}$ precisely since it defines the center of the 3D Gaussian [12]. In this paper, we present MVSplat, a Gaussian-based feed-forward model for novel view synthesis. Unlike pixelSplat [1] that predicts probabilistic depth, we develop an efficient and high-performance multi-view depth estimation model that enables unprojecting predicted depth maps as the Gaussian centers, in parallel with another branch for prediction of other Gaussian parameters ($\alpha_{j}$, $\bm{\Sigma}_{j}$ and $\bm{c}_{j}$). Our full model, illustrated in Fig. 2, is trained end-to-end using only a simple rendering loss for supervision. Next, we discuss key components.
@@ 3.1 Multi-View Depth Estimation
Multi-view feature extraction. To construct the cost volumes, we first extract multi-view image features with a CNN and Transformer architecture [34, 35]. More specifically, a shallow ResNet-like CNN is first used to extract 4×4$\times$4 × downsampled per-view image features, and then we use a multi-view Transformer with self- and cross-attention layers to exchange information between different views. For better efficiency, we use Swin Transformer’s local window attention [15] in our Transformer architecture. When more than two views are available, we perform cross-attention for each view with respect to all the other views, which has exactly the same learnable parameters as the 2-view scenario. After this operation, we obtain cross-view aware Transformer features $\{{\bm{F}}^{i}\}_{i=1}^{K}$ (${\bm{F}}^{i}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times C}$), where $C$ denotes the channel dimension.
Cost volume construction. An important component of our model is the cost volume, which models cross-view feature matching information with respect to different depth candidates via the plane-sweep stereo approach [37, 35]. Note that we construct $K$ cost volumes for $K$ input views to predict $K$ depth maps. Here, we take view $i$’s cost volume construction as an example. Given the near and far depth ranges, we first uniformly sample $D$ depth candidates $\{d_{m}\}_{m=1}^{D}$ in the inverse depth domain and then warp view $j$’s feature ${\bm{F}}^{j}$ to view $i$ with the camera projection matrices ${\bm{P}}^{i}$, ${\bm{P}}^{j}$ and each depth candidate $d_{m}$, to obtain $D$ warped features
@@
To exchange information between cost volumes of different views, we inject three cross-view attention layers at the lowest resolution of the U-Net architecture. The cross-view attention layers are able to accept arbitrary number of views as input since it computes the cross-attention for each view with respect to all the other views, where such an operation does not depend on the number of views. The low-resolution cost volume ${\tilde{\bm{C}}}^{i}$ is finally upsampled to full resolution ${\hat{\bm{C}}}^{i}\in\mathbb{R}^{H\times W\times D}$ with a CNN-based upsampler.
@@
Depth refinement. To further improve the performance, we introduce an additional depth refinement step to enhance the quality of the above predicted depth. The refinement is performed with a very lightweight 2D U-Net, which takes multi-view images, features and current depth predictions as input, and outputs per-view residual depths. The residual depths are then added to the current depth predictions as the final depth outputs. Similar to the U-Net used in the above cost volume refinement, we also introduce cross-view attention layers in the lowest resolution to exchange information across views. More implementation details are presented in the supplementary material Appendix 0.C.
@@ 3.2 Gaussian Parameters Prediction
Covariance $\bm{\Sigma}$ and color $\bm{c}$. We predict these parameters by two convolution layers that take as inputs the concatenated image features, refined cost volume and the original multi-view images. Similar to other 3DGS approaches [12, 1], the covariance matrix $\bm{\Sigma}$ is composed of a scaling matrix and a rotation matrix represented via quaternions, and the color $\bm{c}$ is calculated from the predicted spherical harmonic coefficients.
@@ 4.1 Settings
Datasets. We assess our model on the large-scale RealEstate10K [42] and ACID [14] datasets. RealEstate10K contains real estate videos downloaded from YouTube, which are split into 67,477 training scenes and 7,289 testing scenes, while ACID contains nature scenes captured by aerial drone, split into 11,075 training scenes and 1,972 testing scenes. Both datasets provide estimated camera intrinsic and extrinsic parameters for each frame. Following pixelSpalt [1], we evaluate all methods on three target novel viewpoints for each test scene. Furthermore, to further evaluate the cross-dataset generalization ability, we also directly evaluate on the multi-view DTU [10] dataset, which contains object-centric scenes with camera poses. On DTU dataset, we report results on 16 validation scenes, with 4 novel views for each scene.
@@
Implementation details. Our MVSplat is implemented with PyTorch, along with an off-the-shelf 3DGS render implemented in CUDA. Our multi-view Transformer contains 6 stacked self- and cross-attention layers. We sample 128 depth candidates when constructing the cost volumes in all the experiments. All models are trained on a single A100 GPU for 300,000 iterations with the Adam [13] optimizer. More details are provided in the supplementary material Appendix 0.C. Code and models are available at https://github.com/donydchen/mvsplat.

### v2

# donydchen.github.io/mvsplat

@@ Abstract
We introduce MVSplat, an efficient model that, given sparse multi-view images as input, predicts clean feed-forward 3D Gaussians. To accurately localize the Gaussian centers, we build a cost volume representation via plane sweeping, where the cross-view feature similarities stored in the cost volume can provide valuable geometry cues to the estimation of depth. We also learn other Gaussian primitives’ parameters jointly with the Gaussian centers while only relying on photometric supervision. We demonstrate the importance of the cost volume representation in learning feed-forward Gaussians via extensive experimental evaluations. On the large-scale RealEstate10K and ACID benchmarks, MVSplat achieves state-of-the-art performance with the fastest feed-forward inference speed (22 fps). More impressively, compared to the latest state-of-the-art method pixelSplat, MVSplat uses 10×10$\times$10 × fewer parameters and infers more than 2×2$\times$2 × faster while providing higher appearance and geometry quality as well as better cross-dataset generalization.
@@ 1 Introduction
We consider the problem of 3D scene reconstruction and novel view synthesis from very sparse (i.e., as few as two) images in just one forward pass of a trained model. While remarkable progress has been made using neural scene representations, e.g., Scene Representation Networks (SRN) [34], Neural Radiance Fields (NeRF) [24] and Light Filed Networks (LFN) [33], these methods are still not satisfactory for practical applications due to expensive per-scene optimization [26, 39, 43], high memory cost [3, 18, 44] and slow rendering speed [40, 49].
@@
Very recently, several feed-forward Gaussian Splatting methods have been proposed to explore 3D reconstruction from sparse view images, notably Splatter Image [37] and pixelSplat [1]. Splatter Image regresses pixel-aligned Gaussian parameters using a standard image-to-image architecture, which achieves promising results for single-view object-level 3D reconstruction. However, reconstructing a 3D scene from a single image is inherently ill-posed and ambiguous, posing a significant challenge when applied to a more general and larger scene, which is the key focus of our paper. pixelSplat [1] proposes to regress Gaussian parameters for the binocular reconstruction problem. Specifically, it predicts a probabilistic depth distribution for each input view and then samples depths from that predicted distribution. Even though pixelSplat learns cross-view-aware features with an epipolar Transformer, it is still challenging to predict a reliable probabilistic depth distribution solely from image features, making pixelSplat’s geometry reconstruction of comparably low quality and exhibiting noisy artifacts (see Fig. 1 and Fig. 4). For improved geometry reconstruction results, slow depth finetuning with an additional depth regularization loss is required.
@@
To accurately localize the 3D Gaussian centers, our solution is to build a cost volume representation via plane sweeping [7, 48, 46] in the 3D space. Specifically, the cost volume stores cross-view feature similarities for all potential depth candidates, where the similarities can provide valuable geometry cues to the localization of 3D surfaces (i.e., high similarity more likely indicates a surface point). With our cost volume representation, the task is formulated as learning to perform feature matching to identify the Gaussian centers, unlike the data-driven 3D regression from image features in previous works [37, 1]. Such a formulation reduces the task’s learning difficulty, enabling our method to achieve state-of-the-art performance with lightweight model size and fast speed.
@@
We obtain 3D Gaussian centers by unprojecting the multi-view-consistent depths estimated by our constructed multi-view cost volumes with a 2D network. Additionally, we also predict other Gaussian properties (covariance, opacity, and spherical harmonics coefficients), in parallel with the depths. This enables the rendering of novel view images using the predicted 3D Gaussians with the differentiable splatting operation [19]. Our full model MVSplat is trained end-to-end purely with the photometric loss between rendered and ground truth images.
@@
On the large-scale RealEstate10K [54] and ACID [21] benchmarks, MVSplat achieves state-of-the-art performance with the fastest feed-forward inference speed (22 fps). More impressively, compared to the state-of-the-art pixelSplat [1] (see Fig. 1), MVSplat uses 10×10$\times$10 × fewer parameters and infers more than 2×2$\times$2 × faster while providing higher appearance and geometry quality as well as better cross-dataset generalization. Furthermore, extensive ablation studies and analysis underscore the significance of our feature matching-based cost volume design in enabling highly efficient feed-forward 3D Gaussian Splatting models.
@@ 2 Related Work
Sparse View Scene Reconstruction and Synthesis. The original NeRF and 3DGS methods are both designed for very dense views (e.g., 100) as inputs, which can be tedious to capture for real-world applications. Recently, there have been growing interests [26, 39, 3, 5, 44, 1, 43, 36, 12, 51] in scene reconstruction and synthesis from sparse input views (e.g., 2 or 3). Existing sparse view methods can be broadly classified into two categories: per-scene optimization and cross-scene feed-forward inference methods. Per-scene approaches mainly focus on designing effective regularization terms [26, 39, 50, 8, 43, 11] to better constrain the optimization process. However, they are inherently slow at inference time due to the expensive per-scene gradient back-propagation process. In contrast, feed-forward models [49, 3, 5, 44, 1, 37, 12, 51, 36] learn powerful priors from large-scale datasets, so that 3D reconstruction and view synthesis can be achieved via a single feed-forward inference by taking sparse views as inputs, which makes them significantly faster than those per-scene optimization methods.
@@
Feed-Forward NeRF. Early approaches used NeRF [24] for objects [6, 16, 18, 22, 27, 40, 49] and scenes [6, 10, 3, 5, 44] 3D reconstrution. pixelNeRF [49] pioneered the paradigm of predicting pixel-aligned features from images for radiance field reconstruction. The performance of feed-forward NeRF models progressively improved with the use of feature matching information [3, 5], Transformers [30, 10, 25] and 3D volume representations [3, 44]. The state-of-the-art feed-forward NeRF model MuRF [44] is based on a target view frustum volume and a (2+1)D CNN for radiance field reconstruction, where the 3D volume and CNN need to be constructed and inferred for every target view. This makes MuRF expensive to train, with comparably slow rendering. Most importantly, all existing feed-forward NeRF models suffer from the expensive per-pixel volume sampling in the rendering process.
@@
Feed-Forward 3DGS. 3D Gaussian Splatting [19, 4] avoids NeRF’s expensive volume sampling via a rasterization-based splatting approach, where novel views can be rendered very efficiently from a set of 3D Gaussian primitives. Very recently, a growing number of feed-forward 3DGS models [37, 1, 53, 42, 51, 2, 36] have been proposed to solve the sparse-view-to-3D task. Among them, Splatter Image [37] proposes to regress pixel-aligned Gaussian parameters from a single view with a U-Net model. However, it mainly focuses on object-level reconstruction, while we target the more general scene-level reconstruction. Although its follow-up work Flash3D [36] manages to extend to scene-level reconstruction, its performance on complex scenes is inherently non-satisfactory due to the limited information a single image can provide. In contrast, MVSplat is designed to effectively aggregate information from multi-view input. More similar to our setting, pixelSplat [1] proposes to regress Gaussian parameters from two input views. It demonstrates the importance of cross-view-aware features, learned from the epipolar Transformer. It then directly uses these features to predict probabilistic depth distributions for sampling depths. However, the mapping from features to depth distributions is inherently ambiguous and unreliable, essentially leading to poor geometry reconstruction. In contrast, we learn to predict depth from the feature matching information encoded within a cost volume, which makes it more geometry-aware and leads to a more lightweight model (10×10$\times$10 × fewer parameters and more than 2×2$\times$2 × faster) and significantly better geometries. Another related work GPS-Gaussian [53] proposes a feed-forward Gaussian model for the reconstruction of humans, instead of general scenes. It relies on two rectified stereo images to estimate the disparity, while our method works for general unrectified multi-view posed images. During training, GPS-Gaussian requires ground truth depth for supervision, while our model is trained from RGB images alone.
Figure 2: Overview of MVSplat. Given multiple posed images as input, MVSplat first extracts multi-view image features with a Transformer. Then, the per-view cost volumes using plane sweeping are constructed. The Transformer features and cost volumes are concatenated together as input to a 2D U-Net (with cross-view attention) for cost volume refinement and predicting per-view depth maps. The per-view depth maps are unprojected to 3D and combined using a simple deterministic union operation as the 3D Gaussian centers. The opacity, covariance and color Gaussian parameters are predicted jointly with the depth maps. Finally, novel views are rendered from the predicted 3D Gaussians with the rasterization operation.
@@ 3 Method
To enable high-quality rendering and reconstruction, it is crucial to predict the position $\bm{\mu}_{j}$ precisely since it defines the center of the 3D Gaussian [19]. In this paper, we present MVSplat, a Gaussian-based feed-forward model for novel view synthesis. Unlike pixelSplat [1] that predicts probabilistic depth, we develop an efficient and high-performance multi-view depth estimation model that enables unprojecting predicted depth maps as the Gaussian centers, in parallel with another branch for prediction of other Gaussian parameters ($\alpha_{j}$, $\bm{\Sigma}_{j}$ and $\bm{c}_{j}$). Our full model, illustrated in Fig. 2, is trained end-to-end using only a simple rendering loss for supervision. Next, we discuss the key components.
@@ 3.1 Multi-View Depth Estimation
Multi-view feature extraction. To construct the cost volumes, we first extract multi-view image features with a CNN and Transformer architecture [45, 46]. Specifically, a shallow ResNet-like CNN is first used to extract 4×4$\times$4 × downsampled per-view image features. Then, we use a multi-view Transformer with self- and cross-attention layers to exchange information between different views. For better efficiency, we use Swin Transformer’s local window attention [23] in our Transformer architecture. When more than two views are available, we perform cross-attention for each view with respect to all the other views, which has exactly the same learnable parameters as the 2-view scenario. After this operation, we obtain cross-view aware Transformer features $\{{\bm{F}}^{i}\}_{i=1}^{K}$ (${\bm{F}}^{i}\in\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times C}$), where $C$ denotes the channel dimension.
Cost volume construction. The key component of our model is the cost volume, which models cross-view feature matching information with respect to different depth candidates via the plane-sweep stereo approach [7, 48, 46]. Note that we construct $K$ cost volumes for $K$ input views to predict $K$ depth maps. Here, we take view $i$’s cost volume construction as an example. Given the near and far depth ranges, we first uniformly sample $D$ depth candidates $\{d_{m}\}_{m=1}^{D}$ in the inverse depth domain and then warp view $j$’s feature ${\bm{F}}^{j}$ to view $i$ with the camera projection matrices ${\bm{P}}^{i}$, ${\bm{P}}^{j}$ and each depth candidate $d_{m}$, to obtain $D$ warped features
@@
To exchange information between cost volumes of different views, we inject three cross-view attention layers at the lowest resolution of the U-Net architecture. This design ensures that the model can accept an arbitrary number of views as input since it computes the cross-attention for each view with respect to all the other views, where such an operation does not depend on the number of views. The low-resolution cost volume ${\tilde{\bm{C}}}^{i}$ is finally upsampled to full resolution ${\hat{\bm{C}}}^{i}\in\mathbb{R}^{H\times W\times D}$ with a CNN-based upsampler.
@@
Depth refinement. To further improve the performance, we introduce an additional depth refinement step to enhance the quality of the predicted depth. The refinement is performed with a very lightweight 2D U-Net, which takes multi-view images, features, and current depth predictions as input, and outputs per-view residual depths. The residual depths are then added to the current depth predictions as the final depth outputs. Similar to the U-Net used in the above cost volume refinement, we also introduce cross-view attention layers in the lowest resolution to exchange information across views. More implementation details are presented in the supplementary material Appendix 0.C.
@@ 3.2 Gaussian Parameters Prediction
Covariance $\bm{\Sigma}$ and color $\bm{c}$. We predict these parameters using two convolution layers that take as inputs the concatenated image features, refined cost volume, and original multi-view images. Similar to other 3DGS approaches [19, 1], the covariance matrix $\bm{\Sigma}=R(\theta)^{\mathsf{T}}\operatorname{diag}(s)R(\theta)$ is composed of a scaling matrix $s$ and a rotation matrix $R(\theta)$ represented via quaternions, and the color $\bm{c}$ is calculated from the predicted spherical harmonic coefficients.
@@ 4.1 Settings
Datasets. We assess our model on the large-scale RealEstate10K [54] and ACID [21] datasets. RealEstate10K contains real estate videos downloaded from YouTube, which are split into 67,477 training scenes and 7,289 testing scenes, while ACID contains nature scenes captured by aerial drones, which are split into 11,075 training scenes and 1,972 testing scenes. Both datasets provide estimated camera intrinsic and extrinsic parameters for each frame. Following pixelSpalt [1], we evaluate all methods on three target novel viewpoints for each test scene. Furthermore, to further assess the cross-dataset generalization ability, we also directly evaluate all models on the multi-view DTU [17] dataset, which contains object-centric scenes with camera poses. On the DTU dataset, we report results on 16 validation scenes, with 4 novel views for each scene.
@@
Implementation details. MVSplat is implemented with PyTorch, along with an off-the-shelf 3DGS render implemented in CUDA. Our multi-view Transformer contains 6 stacked self- and cross-attention layers. We sample 128 depth candidates when constructing the cost volumes in all the experiments. All models are trained on a single A100 GPU for 300,000 iterations with the Adam [20] optimizer. More details are provided in the supplementary material Appendix 0.C. Code and models are available at https://github.com/donydchen/mvsplat.